# Acquiring Bidirectionality via Large and Small Language Models

### v1

@@ Acknowledgments
Foward-LM only - - B-PER - B-ORG - - B-ORG - - - -
Proposal - - B-ORG - B-ORG - B-MISC B-ORG - - - -
@@
Table 4: An examples of GPT-2 (base) in phrases where entities are conjoined by “and”. For the prediction corresponding to Lotte, the forward-LM struggles to infer the entity type correctly, but proposal can estimate correctly. This can be explained by the difference between considering only the context from the beginning: “Note - Lotte” or entire context, cause the improvement.
@@
Appendix A Classification Layer with Transformer Block
We use a one-layer Transformer layer: $Transformer(\cdot):768\mapsto 768$, which the same as BERT-base configuration. The final distribution is calculated by $\mathbf{p}=\mathrm{softmax}(\mathbf{W}_{2}^{trans}Transformer(\mathbf{W}_{1}^{trans}\overrightarrow{\mathbf{h}}_{i}))\in\mathbb{R}^{c}$, where $\mathbf{W}_{1}^{trans}\in\mathbb{R}^{768\times d}$ and $\mathbf{W}_{2}^{trans}\in\mathbb{R}^{c\times 768}$ are dense layers to align the input/output dimensions.
Appendix B Case studies
Table 3 and 4 show the actual outputs of both forward LM-only setting and proposal setting.

### v2

@@ Acknowledgments
Forward-LM only - - B-PER - B-ORG - - B-ORG - - - -
Proposed - - B-ORG - B-ORG - B-MISC B-ORG - - - -
@@
Table 2: An example of GPT-2 (base) in phrases where entities are conjoined by “and.” For the prediction corresponding to Lotte, the forward-LM struggles to infer the entity type correctly, but proposal can estimate correctly. This can be explained by the difference between considering only the context from the beginning: “Note - Lotte” or entire context, cause the improvement.
@@
Appendix A Another Example in Case Study
Table 2 shows an example where entities are concatenated with “and.” The forward-LM only setting struggles with identifying the type of first entity, for “Lotte.” In contrast, the proposed method correctly identified the entity type of it by using other organization names such as “Hyundai.”